# Energy-Efficiency and Spectral-Efficiency Trade-off in Distributed Massive-MIMO Networks

### v1

@@ I Introduction
Distributed massive-MIMO (D-mMIMO) combines the advantages of massive-MIMO and the cell-free architecture to improve both SE and EE. D-mMIMO systems provide uniform service quality over a large coverage area by distributing antennas across the region, significantly enhancing both EE and SE compared to conventional cellular networks [1].
@@ II System Model
We consider a scenario where $T$ single antenna UEs and $M$ multi-antenna APs operate within a defined coverage area. Each AP is equipped with $A$ antennas, and the relationship $T\ll MA$ indicates a high antenna-to-user ratio, facilitating substantial spatial multiplexing gains. The APs and the UEs are uniformly distributed across the coverage area, ensuring that every UE is within the service range of multiple APs. All APs can serve all UEs simultaneously using the same time and frequency resources. Connectivity between the APs and a central processing unit (CPU) is ensured via a front-haul link, which is critical for coordinating signal processing tasks and distributing user data and control information efficiently across the network. This setup uses Time-Division Duplexing (TDD) mode for communication. Channel estimation is conducted during the uplink phase where the UEs transmit orthogonal pilot signals, and the APs estimate the uplink channel vectors. The system employs a block fading model, characterized by a coherence block of length $L_{c}$. Within each coherence block, a portion defined by $L_{p}$ is dedicated to pilot transmission for channel estimation. The small-scale fading component, denoted by $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, follows a Rayleigh distribution. Each component of $\textbf{h}_{mt}$ is assumed to be an independent and identically distributed (i.i.d.) complex Gaussian random variable with zero mean and unit variance. Large-scale fading coefficient (LSFC), denoted by $\beta_{mt}$, accounts for path-loss and shadowing effects. The overall channel vector $\textbf{g}_{mt}\in\mathbb{C}^{A\times 1}$ is given by the product $\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$.
@@ IV The SCA based Energy Efficiency maximization
Step 1: In the objective function (5a), the numerator and denominator both contain the sum SE term, so instead of maximizing objective function (5a), we maximizes new objective function $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}}$. Then the new problem becomes:
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+\bar{P}_{\text{T}}},$ (6a)
@@
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})}{P^{\text{fix}}_{\text{T}}+\bar{P}_{\text{T}}},$ (12a)
@@ VI Conclusion and Future Work
A compelling future direction would involve drawing detailed comparisons of the EE-SE trade-offs across different network architectures, including small-cell networks, traditional mMIMO and the D-mMIMO.

### v2

@@ I Introduction
Distributed massive-MIMO (D-mMIMO) combines the advantages of massive-MIMO and the cell-free architecture to improve performance of the network [1]. D-mMIMO systems provide uniform service quality over a large coverage area by distributing antennas across the region, significantly enhancing both EE and SE compared to conventional cellular networks [2].
@@ II System Model
We consider a scenario where $T$ single antenna UEs and $M$ multi-antenna APs operate within a defined coverage area. Each AP is equipped with $A$ antennas, and the relationship $T\ll MA$ indicates a high antenna-to-user ratio, facilitating substantial spatial multiplexing gains. The APs and the UEs are uniformly distributed across the coverage area, ensuring that every UE is within the service range of multiple APs. All APs can serve all UEs simultaneously using the same time and frequency resources. Connectivity between the APs and the central processing unit (CPU) is ensured via a front-haul link, which is critical for coordinating signal processing tasks and distributing user data and control information efficiently across the network as shown in Fig. 1. To ensure network scalability, a user-centric approach has been adopted, where each UE is served by a subset of APs, as depicted in Fig. 1.
Figure 1: A typical network model. Here each oval identifies the UE(s) and the APs that serve it(them).
This setup uses Time-Division Duplexing (TDD) mode for communication. Channel estimation is conducted during the uplink phase where the UEs transmit orthogonal pilot signals, and the APs estimate the uplink channel vectors. The system employs a block fading model, characterized by a coherence block of length $L_{c}$. Within each coherence block, a portion defined by $L_{p}$ is dedicated to pilot transmission for channel estimation. The small-scale fading component, denoted by $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, follows a Rayleigh distribution. Each component of $\textbf{h}_{mt}$ is assumed to be an independent and identically distributed (i.i.d.) complex Gaussian random variable with zero mean and unit variance. Large-scale fading coefficient (LSFC), denoted by $\beta_{mt}$, accounts for path-loss and shadowing effects. The overall channel vector $\textbf{g}_{mt}\in\mathbb{C}^{A\times 1}$ is given by the product $\textbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$.
@@ IV The SCA based Energy Efficiency maximization
Step 1: In the objective function (5a), the numerator and denominator both contain the sum SE term, so instead of maximizing objective function (5a), we maximize new objective function $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}}$. Then the new problem becomes:
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}},$ (6a)
@@
$\underset{\bm{\eta^{u}},\textbf{D}}{\max~{}}$ $\frac{wB\sum\limits_{t=1}^{T}\log_{2}(1+\Gamma^{*}_{t})}{P^{\text{fix}}_{\text{T}}+{P}^{c}_{\text{T}}},$ (12a)
@@ VI Conclusion and Future Work
A compelling future direction would involve drawing detailed comparisons of the EE-SE trade-offs across different network architectures, including small-cell networks, traditional mMIMO and the D-mMIMO. Additionally, the influence of modulation and coding schemes, which is not considered in this study, presents an interesting direction for further exploration.